A meter stick balances horizontally on a knife-edge at the 50.0 cm mark. With two 5.34 g coins stacked over the 17.9 cm mark, the stick is found to balance at the 42.0 cm mark. What is the mass of the meter stick?
Answer #1

Let m be the mass of the meter stick. We will take moments about the pivot (42 cm mark). The distance of the stacked coins from the pivot is (42-17.9) = 24.1 cm on the left-hand side. The center of mass of the stick is (50-42) = 8 cm away from the pivot on the right-hand side.

We have, (2*5.34)*g * 24.1 = mg*8

=> m = 32.17 g (ans)